OpenStudy (anonymous):

A student is attempting to solve a multi-step equation. Sample mathematical work is shown below. Which statement best applies to the sample mathematical work? Given the equation 2003-05-05-00-00_files/i0300000.jpg, I must solve for m. I first subtract 8 from both sides, which gives me 2003-05-05-00-00_files/i0300001.jpg. I then multiply both sides by –2, which gives me the final answer of 2003-05-05-00-00_files/i0300002.jpg. A. The student incorrectly subtracted 8 from both sides. B. The student incorrectly multiplied both sides by –2. C. The final answer is not reduced or simp

OpenStudy (anonymous):

whoops

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

A student is attempting to solve a multi-step equation. Sample mathematical work is shown below. Which statement best applies to the sample mathematical work? Given the equation , I must solve for m. I first subtract 8 from both sides, which gives me \[-m/2\] = -25/3. I then multiply both sides by –2, which gives me the final answer of m=50/3 A. The student incorrectly subtracted 8 from both sides. B. The student incorrectly multiplied both sides by –2. C. The final answer is not reduced or simplified. D. The mathematical work shown is correct